OpenStudy (phoenixfire):

A free electron in an oscillating field experiences a force \(\mathbf{F}=-e\mathbf {E}(t)\), where \(\mathbf {E}(t)=\mathbf {E_0}sin(wt)\) and \(\mathbf {E_0}=(E_0,0,0)\) <- completely in the x direction. if x(t) is the x-coordinate of the electron and we assume \(x(0)=\frac{dx}{dt}(0)=0\) How do we find the equation of motion along the 'x' axis from this information?

OpenStudy (anonymous):

\[m \frac{d^2x}{dt^2} + eE_oSin(\omega t) = 0\] solve this differential equatinon :D

OpenStudy (phoenixfire):

@Mashy how did you come up with that differential?

OpenStudy (phoenixfire):

\(F=ma=-eE(t)\) \(m\frac{d^2 x}{dt^2}=-eE(t)\) Is this what you did?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yes yes.. sorry.. i was in a hurry and so i forgot to mention :D